OpenStudy (anonymous):

Find the equation of the tangent line to the curve at the given point. \\ y=1+2x-x^3, (1,2) An explanation would be appreciated.

OpenStudy (badhi):

If the gradient and a point is given, we can find the equation of the line. A point on the tangent line is already given. All you have to find now is the gradient. Apply differentiation to the given curve for that.

OpenStudy (tkhunny):

I always puzzle over this kind of presentation. Were you given this problem with no explanation whatsoever? Something is seriously wrong with that! I blame your teacher or instructor if you have one - whoever wrote the curriculum if you don't. Anyway... y=1+2x-x^3, (1,2) First, make sure the point is ON the curve. 2 = 1+2(1)-(1)^3 = 1 + 2 - 1 = 2 -- Okay, we're good. We need the equation of a line passing through (1,2) y - k = m(x-h) -- In Point-Slope Form We know the point. y - 2 = m(x-1) Where shall we find the slope at x = 1? The answer for now is the 1st Derivative. y' = 2 -3x^2 Evaluating at x = 1, y'(1) = 2 - 3 = -1 <== This is the required slope. y - 2 = -(x-1) -- Done!

OpenStudy (08surya):